OpenStudy (anonymous):

I'm trying to find the total current, the formula is I=V/R If for V I have 12 and for R I have a fraction(Ex: 10/8) How would I make the formula work?? Would I turn the fraction 10/8 into a decimal?

OpenStudy (anonymous):

You could express the fraction as a decimal, then divide 12 by that value. Dividing by a fraction is the same as multiplying by the reciprocal of the fraction, so you could also multiply 12 by the reciprocal of 10/8.

OpenStudy (anonymous):

You said I can express the fraction as a decimal, you mean 10/8 right?

OpenStudy (anonymous):

If you have 12 / (10 / 8) that is the same as having (12 / 10) * 8 = 1.2 * 8 = 9.6

OpenStudy (anonymous):